For a chemical reaction, temperature is a function of time. For Denver Poweroxide this function is T = -t^2 + 7t + 8 (T is degrees Fahrenheit and t is seconds). How long will it take for the temperature to reach zero? Solve by graphing.
I got: 1 second. Is this correct?
Thanks~

Answer by stanbon(75887)

T = -t^2 + 7t + 8
T=-0
-t^2+2t+8=0
Multiply thru by -1 to get:
t^2-2t-8=0
Factor to get:
(t-4)(t+2)=0
t=4 seconds is the only positive answer.
=========
Cheers,
Stan H.